### v1

# A Note on Approximating Weighted Nash Social Welfare with Additive Valuations

Abstract
We give the first $O(1)$-approximation for the weighted Nash Social Welfare problem with additive valuations. The approximation ratio we obtain is $e^{1/e}+\epsilon\approx 1.445+\epsilon$, which matches the best known approximation ratio for the unweighted case [3].
Both our algorithm and analysis are simple. We solve a natural configuration LP for the problem, and obtain the allocation of items to agents using a randomized version of the Shmoys-Tardos rounding algorithm developed for unrelated machine scheduling problems [30]. In the analysis, we show that the approximation ratio of the algorithm is at most the worst gap between the Nash social welfare of the optimum allocation and that of an EF1 allocation, for an unweighted Nash Social Welfare instance with identical additive valuations. This was shown to be at most $e^{1/e}\approx 1.445$ by Barman et al. [3], leading to our approximation ratio.
1 Introduction
In the weighted (or asymmetric) Nash Social Welfare problem with additive valuations, we are given a set ${\mathcal{A}}$ of $n$ agents, and a set ${\mathcal{G}}$ of $m$ indivisible items. Every agent $i\in{\mathcal{A}}$ has a weight $w_{i}\geq 0$ such that $\sum_{i\in{\mathcal{A}}}w_{i}=1$. There is a value $v_{ij}\in{\mathbb{R}}_{\geq 0}$ for every $i\in{\mathcal{A}}$ and $j\in{\mathcal{G}}$. The goal of the problem is to find an allocation $\sigma:{\mathcal{G}}\to{\mathcal{A}}$ of items to agents so as to maximize the following weighted Nash social welfare of $\sigma$:
$\prod_{i\in{\mathcal{A}}}\left(\sum_{j\in\sigma^{-1}(i)}v_{ij}\right)^{w_{i}}.$
In the case where all $w_{i}$’s are equal to $\frac{1}{n}$, we call the problem the unweighted (or symmetric) Nash Social Welfare problem.
Allocating resources in a fair and efficient manner among multiple agents is a fundamental problem in computer science, game theory, and economics, with applications across diverse domains [19, 33, 4, 28, 25, 2, 29, 5]. The weighted Nash social welfare function is a notable objective that balances efficiency and fairness. The unweighted (or symmetric) objective was independently proposed by different communities [26, 20, 32], and later the study has been extended to the weighted case [16, 18]. Since then it has been used in a wide range of applications, including bargaining theory [21, 7, 31], water allocation [17, 10], and climate agreements [34].
The unweighted Nash Social Welfare problem with additive valuations is proved to be NP-hard by Nguyen et al. [27], and APX-hard by Lee [22]. Later the hardness of approximation was improved to $\sqrt{8/7}\approx 1.069$ by Garg et al. [12], via a reduction from Max-E3-Lin-2.
On the positive side, Cole and Gkatzelis [9] gave a $(2e^{1/e}+\epsilon\approx 2.889+\epsilon)$-approximation using a market equilibrium with some spending restrictions. The ratio was improved by Cole et al. [8] to $2$ using a tight analysis, and by Anari et al. [1] to $e$ via a connection of the problem to real stable polynomials. Both papers formulated some convex program (CP) relaxations for the problem. In particular, [8] showed that the optimum solution to their CP corresponds to the spending-restricted market equilibrium defined in [9]. The state-of-the-art result for the problem is a combinatorial $(e^{1/e}+\epsilon\approx 1.45+\epsilon)$-approximation algorithm due to Barman et al.[3]. They showed that when all the valuations of agents are identical, any allocation that is envy-free up to one item (EF1) is $e^{1/e}$-approximate. Their approximation result then follows from a connection between the non-identical and identical valuation settings they established.
All the results discussed above are for the unweighted case. For the weighted case with agent weights $w\in[0,1]^{\mathcal{A}},|w|_{1}=1$, Brown et al. [6] presented a $5\cdot\exp(2\cdot D_{\text{KL}}\Big{(}w||\frac{\vec{1}}{n})\Big{)}=5\cdot\exp(2\log n+2\sum_{i\in{\mathcal{A}}}w_{i}\log w_{i})$ approximation algorithm, where $D_{\text{KL}}$ denotes the KL divergence of two distributions. This is the first work that studies the weighted version for the additive valuation case. Prior to this work, there is an $O(nw_{\max})=O(n\max_{i\in{\mathcal{A}}}w_{i})$-approximation for the more general submodular valuation case [13], which we discuss soon. Brown et al. [6] showed that the two CPs from [8] and [1] are equivalent, and their result is based on the CP from [8], generalized to the weighted setting.
The additive valuation setting is a special case of the submodular valuation setting, which is another important setting studied in the literature. In this setting, instead of a $v_{ij}$ value for every $ij$ pair, we are given a monotone submodular function $v_{i}:2^{{\mathcal{G}}}\to{\mathbb{R}}_{\geq 0}$ for every agent $i\in{\mathcal{A}}$. Our goal is to find an allocation $\sigma:{\mathcal{G}}\to{\mathcal{A}}$ so as to maximize $\prod_{i\in{\mathcal{A}}}\Big{(}v_{i}(\sigma^{-1}(i))\Big{)}^{w_{i}}$. A bulk of the previous work has focused on the unweighted case; that is, $w_{i}=\frac{1}{n}$ for all $i\in{\mathcal{A}}$. For this case, Garg et al. [15] proved a hardness of $e/(e-1)\approx 1.5819$ using a reduction from Max-3-Coloring; this is better than the $1.069$ hardness for the additive valuation case.
On the positive side, Li and Vondrak [24] extended the techniques of Anari et al. [1], to obtain an $e^{3}/(e-1)^{2}$-approximation algorithm for the unweighted Nash Social Welfare problem for a large family of submodular valuations, including coverage functions and linear combinations of matroid rank functions. Later, Garg et al. [14] considered a family of submodular functions called Rado functions, and gave an $O(1)$-approximation for this family using the matching theory and convex program techniques. Li and Vondrak [23] developped the first $O(1)$-approximation for general submodular functions, with an approximation ratio of $380$. Recently, Garg et al. [13] presented an elegant $4$-approximation local search algorithm for the problem, which is the current best approximation result for the problem. All the results discussed above are for the unweighted case. For the weighted case, Garg et al. [13] gave an $O(nw_{\max})$-approximation, where $n_{\max}=\max_{i\in{\mathcal{A}}}w_{i}$. Whether the weighted Nash Social Welfare problem with submodular valuations admits a constant approximation is a big open problem.
Recently, the problem has been studied in an even more general setting, namely, the subadditive valuation setting. Dobzinski et al. [11] gives an $O(1)$-approximation for the unweighted Nash Social Welfare problem in this setting under the demand oracle model.
1.1 Our Result and Techniques
In this note, we give the first $O(1)$-approximation algorithm for the weighted Nash Social Welfare problem with additive valuations:
Theorem 1.1.
For any $\epsilon>0$, there is a randomized $(e^{1/e}+\epsilon\approx 1.445+\epsilon)$-approximation algorithm for the weighted Nash Social Welfare problem with additive valuations, with running time polynomial in the size of the input and $\frac{1}{\epsilon}$.
Our approximation ratio of $e^{1/e}+\epsilon$ matches the best ratio for the unweighted case due to Barman et al. [3]. In contrast, the ratio given by Brown et al. [6] is $5\cdot\exp(2\cdot D_{\text{KL}}(w||\frac{\vec{1}}{n}))$, which could be polynomial in $n$.
Our algorithm is based on a natural configuration LP for the problem, which has not been studied before to the best of our knowledge. The configuration LP contains a $y_{i,S}$ variable for every agent $i$ and subset $S$ of items, indicating if the set of items $i$ gets is $S$ or not. We show that the configuration LP can be solved in polynomial time to any precision, despite having exponential number of variables. Once we obtain the LP solution, we define $x_{ij}$ for every $i\in{\mathcal{A}}$ and $j\in{\mathcal{G}}$ to be the fraction of $j$ assigned to $i$.
We use a randomized version of the Shmoys-Tardos rounding algorithm [30] developed for unrelated machine scheduling problems, to round $x$ into an integral solution. For every agent $i$, we break the fractional items assigned to $i$ into groups from the most valuable to the least, each containing 1 fractional item. The rounding algorithm maintains marginal probabilities, and the requirement that $i$ gets exactly one item from each group (except for the last one, from which $i$ gets at most one item). In the analysis for each agent $i$, we construct an instance of the unweighted Nash Social Welfare problem with identical additive valuations, that involves many copies of the agent $i$, along with two alloations ${\mathcal{S}}$ and ${\mathcal{S}}^{\prime}$ to the instance. ${\mathcal{S}}$ corresponds to the LP solution, and ${\mathcal{S}}^{\prime}$ corresponds to the randomized solution given by the rounding algorithm. Thanks to the condition that every group contains one item, the solution ${\mathcal{S}}^{\prime}$ is envy-free up to one item (EF1). Using the result of [3] about EF1 allocations, we show that the Nash social welfare of ${\mathcal{S}}^{\prime}$ is at least $e^{-1/e}$ times that of ${\mathcal{S}}$, which eventually leads to our $(e^{1/e}+\epsilon)$-approximation.
We believe the configuration LP could be used in many other settings. We leave as an immediate open problem whether it can give an $O(1)$-approximation for the weighted Nash Social Welfare problem with submodular valuations.
2 $(e^{1/e}+\epsilon)$-Approximation Using Configuration LP
We describe the configuration LP in Section 2.1 and the rounding algorithm in Section 2.2. The analysis is given in Section 2.3.
2.1 The Configuration LP
For convenience, for any value function $v:{\mathcal{G}}\to{\mathbb{R}}_{\geq 0}$, we define $v(S):=\sum_{j\in S}v_{j}$ for every $S\subseteq{\mathcal{G}}$ to be the total value of items in $S$ according to the value funciton $v$. In the integer program correspondent to the configuration LP, for every $i\in\cal A$ and $S\subseteq{\mathcal{G}}$, we have a variable $y_{i,S}\in\{0,1\}$ indicating if the set of items assigned to $i$ is $S$ or not. We relax the integer constraint to obtain the following configuration LP:
[…]
$\sum_{i\in{\mathcal{A}},S\ni j}y_{i,S}$ $\leq 1$ $\forall j\in{\mathcal{G}}$ (1)
$\sum_{S\subseteq{\mathcal{G}}}y_{i,S}$ $=1$ $\forall i\in{\mathcal{A}}$ (2)
$y_{i,S}$ $\geq 0$ $\forall i\in{\mathcal{A}},S\subseteq{\mathcal{G}}$ (3)
It is convenient for us to consider the natural logarithm of the Nash social welfare function as the objective, which is $\sum_{i\in{\mathcal{A}}}w_{i}\cdot\ln v_{i}(\sigma^{-1}(i))$. This leads to the objective in (Conf-LP). (1) requires that every item $j$ is assigned to at most one agent, and (2) requires that every agent $i$ is assigned one set of items.
The configuration LP has exponential number of variables, but it can be solved within an additive error of $\ln(1+\epsilon)$ for any $\epsilon>0$, in time polynomial in the size of the instance and $\frac{1}{\epsilon}$. We defer the details to Appendix A. Notice that we are considering the logarithm of Nash social welfare, and the typical $(1+\epsilon)$-multiplicative factor becomes an additive error of $\ln(1+\epsilon)$.
2.2 The Rounding Algorithm
From now on, we assume we have obtained a vector $y$ from solving the LP, described using a list of non-zero coordinates; the value of $y$ to (Conf-LP) is at least the optimum value minus $\ln(1+\epsilon)$. We can assume (1) holds with equalities: $\sum_{i\in{\mathcal{A}},S\ni j}y_{i,S}=1$ for every $j\in{\mathcal{G}}$. Then we let $x_{ij}=\sum_{S\ni j}y_{i,S}$ for every $i\in{\mathcal{A}}$ and $j\in{\mathcal{G}}$. So $\sum_{i\in{\mathcal{A}}}x_{ij}=1$ for every $j\in{\mathcal{G}}$.
In this paragraph, we fix an agent $i$ and break the fractional items assigned to $i$ into a set $G_{i}$ of groups, each containing 1 fractional item. They are created in non-increasing order of values, as in the Shmoys-Tardos algorithm for unrelated machine scheduling problems. That is, the first group contains the 1 fractional most valuable items assigned to $i$, the second group contains the 1 fractional most valuable items assigned to $i$ after removing the first group, and so on. Formally, we sort the items in ${\mathcal{G}}$ in non-increasing order of $v_{ij}$ values, breaking ties arbitrarily. Let $p_{i}={\lceil\sum_{j\in{\mathcal{G}}}x_{ij}\rceil}$. Then we can find vectors $g^{1},g^{2},\cdots,g^{p_{i}}\in[0,1]^{{\mathcal{G}}}$ satisfying the following properties:
  1. (P1)
    For every $t\in[1,p_{i}-1]$, we have $|g^{t}|_{1}=1$; for $t=p_{i}$, we have $|g^{t}|_{1}=\sum_{j\in{\mathcal{G}}}x_{ij}-(p_{i}-1)\in(0,1]$.
  2. (P2)
    $\sum_{t=1}^{p_{i}}g^{t}_{j}=x_{ij}$ for every $j\in{\mathcal{G}}$.
  3. (P3)
    For every $1\leq t<t^{\prime}\leq p_{i}$, and two items $j,j^{\prime}$ such that $j$ appears before $j^{\prime}$ in the ordering, it can not happen that $g^{t}_{j^{\prime}}>0$ and $g^{t^{\prime}}_{j}>0$.
It is easy to see that $g^{1},g^{2},\cdots,g^{p_{i}}$ are uniquely decided by the three conditions. We say each $g^{t}$ is a group. Let $G_{i}=\{g^{1},g^{2},\cdots,g^{p_{i}}\}$ be the set of all groups constructed for this agent $i$.
Now we take all agents $i$ into consideration and let $G=\uplus_{i\in{\mathcal{A}}}G_{i}$ be the set of all groups constructed. [^1] It is possible that two groups from different sets $G_{i}$ and $G_{i^{\prime}}$ have the same vector representation. So we treat $G$ as a multiset and we assume we know which set $G_{i}$ each group $g\in G$ belongs to. The representations of groups give a fractional matching between the groups $G$ and items ${\mathcal{G}}$: an item $j$ is matched to a group $g\in[0,1]^{{\mathcal{G}}}$ with a fraction of $g_{j}$. Then each item is matched to an extent of 1, and every group $g$ is matched to an extent of $|g|_{1}$. So a group is matched to an extent of 1 if it is not the last group for an agent, and at most 1 otherwise. Therefore, we can efficiently output a randomized (partial-)matching between the groups $G$ and items ${\mathcal{G}}$ so that the marginal probabilities are maintained:
  1. ($\star$)
    For every group $g\in G$ and item $j\in{\mathcal{G}}$, we have $\Pr[j\text{ is matched to }g]=g_{j}$.
($\star$) implies that an item $j\in{\mathcal{G}}$ is matched with probability 1. If a group $g$ has $|g|_{1}=1$, then it is matched with probability 1.
The matching naturally gives us an allocation of items to agents: If an item $j\in{\mathcal{G}}$ is matched to some group $g\in G_{i}$, then we assign $j$ to $i$. By ($\star$) we know that the probability that $j$ is assigned to $i$ is precisely $x_{ij}$. Let $S_{i}$ be the set of items assigned to $i$ in the algorithm; notice that it is random. This finishes the description of the randomized rounding algorithm.
2.3 The Analysis
To analyze our rounding algorithm, we first formally define an EF1 allocation.
Definition 2.1.
Given an instance of the unweighted Nash Social Welfare problem with agents ${\mathcal{A}}$, items ${\mathcal{G}}$, and identical additive valuation $v:{\mathcal{G}}\to{\mathbb{R}}_{\geq 0}$ for all agents, an allocation $\sigma:{\mathcal{G}}\to{\mathcal{A}}$ is said to be envy-free up to one item (EF1), if for every two distinct agents $i,i^{\prime}$ with $\sigma^{-1}(i^{\prime})\neq\emptyset$, there exists some $j\in\sigma^{-1}(i^{\prime})$, such that $v(\sigma^{-1}(i^{\prime})\setminus j)\leq v(\sigma^{-1}(i))$.
We use the following result from [3]:
Theorem 2.2 ([3]).
For the unweighted Nash Social Welfare problem with identical additive valuations, any EF1-allocation is an $e^{1/e}$-approximate solution.
With the theorem, we prove the following key lemma:
Lemma 2.3.
For every $i\in{\mathcal{A}}$, we have
$\operatorname{\mathbb{E}}\big{[}\ln v_{i}(S_{i})\big{]}\geq\sum_{S\subseteq{\mathcal{G}}}y_{i,S}\cdot\ln v_{i}(S)\ -\ \frac{1}{e}.$
Proof.
Throughout the proof, we fix the agent $i$. Let $\Delta>0$ be an integer, so that every $y_{i,S}$ is an integer multiply of $1/\Delta$, and the probability that $S_{i}=S$ for any $S$ is also an integer multiply of $1/\Delta$ [^2] We can assume all $y_{i,S}$ values are rational numbers. Under this condition, it is easy to guarantee that the probabilities are rational numbers. We consider an instance of the unweighted Nash Social Welfare problem with identical additive valuations. In the instance, there are $\Delta$ copies of the agent $i$, and $\Delta x_{ij}$ copies of every item $j\in{\mathcal{G}}$; so all the agents are identical. The $y=(y_{i,S})_{S\subseteq{\mathcal{G}}}$ vector gives us an allocation ${\mathcal{S}}$ to the instance: For every $S\subseteq{\mathcal{G}}$, there are exactly $\Delta y_{i,S}$ agents who get a copy of $S$. Notice that this is a valid solution, as $\sum_{S}y_{i,S}=1$ and $\sum_{S\ni j}y_{i,S}=x_{ij}$ for every item $j$.
The Nash Social Welfare of the allocation ${\mathcal{S}}$ is
$\left(\prod_{S\subseteq{\mathcal{G}}}v_{i}(S)^{\Delta y_{i,S}}\right)^{1/\Delta}=\prod_{S\subseteq{\mathcal{G}}}v_{i}(S)^{y_{i,S}}.$
The distribution for $S_{i}$ also corresponds to an allocation ${\mathcal{S}}^{\prime}$ of items to agents: For every $S\subseteq{\mathcal{G}}$, there are $\Delta\cdot\Pr[S_{i}=S]$ agents who get a copy of $S$. Again, this is a valid solution as $\sum_{S}\Pr[S_{i}=S]=1$ and $\sum_{S\ni j}\Pr[S_{i}=S]=\operatorname{\mathbb{E}}[S_{i}\ni j]=x_{ij}$.
The Nash Social Welfare of the allocation ${\mathcal{S}}^{\prime}$ is
$\left(\prod_{S\subseteq{\mathcal{G}}}v_{i}(S)^{\Delta\Pr[S_{i}=S]}\right)^{1/\Delta}=\prod_{S\subseteq{\mathcal{G}}}v_{i}(S)^{\Pr[S_{i}=S]}.$
A crucial property for the solution ${\mathcal{S}}^{\prime}$ is that it is EF1. Indeed, if $\Pr[S_{i}=S]>0$ for some $S$, then $S$ contains exactly one item from each group in $G_{i}$ except for the last one, from which $S$ contains at most one item. Also, the items in the groups $G_{i}$ are sorted by (P3). So if there are two sets $S$ and $S^{\prime}$ in the support of the distribution for $S_{i}$, and we remove the most valuable item from $S^{\prime}$, then $S$ beats $S^{\prime}$ item by item.
Therefore, by Theorem 2.2, we know that the Nash Social Welfare of ${\mathcal{S}}^{\prime}$ is at least $e^{-1/e}$ times that of the optimum allocation for the instance, which is at least that of ${\mathcal{S}}$. That is,
$\prod_{S\subseteq{\mathcal{G}}}v_{i}(S)^{\Pr[S_{i}=S]}\geq e^{-1/e}\cdot\prod_{S\subseteq{\mathcal{G}}}v_{i}(S)^{y_{i,S}}.$
Taking logarithm on both sides gives the lemma. ∎
Applying the lemma for every $i\in{\mathcal{A}}$ and using linearity of expectation, we have
$\operatorname{\mathbb{E}}\left[\sum_{i\in{\mathcal{A}}}w_{i}\cdot\ln v_{i}(S_{i})\right]\geq\sum_{i\in S,S\subseteq{\mathcal{G}}}w_{i}\cdot y_{i,S}\cdot\ln v_{i}(S)-\frac{1}{e}.$
We used that $\sum_{i\in{\mathcal{A}}}w_{i}=1$.
By the convexity of exponential function, we have
$\operatorname{\mathbb{E}}\left[\prod_{i\in{\mathcal{A}}}v_{i}(S_{i})^{w_{i}}\right]\geq e^{-1/e}\cdot\exp\left(\sum_{i\in S,S\subseteq{\mathcal{G}}}w_{i}\cdot y_{i,S}\cdot\ln v_{i}(S_{i})\right)\geq e^{-1/e}\cdot\frac{\text{opt}}{1+\epsilon},$
where opt is the Nash Social Welfare of the optimum allocation, and the second inequality used that the value of our solution $y$ to (Conf-LP) is at least its optimum value minus $\ln(1+\epsilon)$. By scaling $\epsilon$ down by an absolute constant at the beginning, we can make the right side to be at least $\frac{\text{opt}}{e^{1/e}+\epsilon}$. This finishes the proof of Theorem 1.1.
Finally, we briefly discuss how to derandomize the rounding algorithm. We round the solution to the configuration LP in iterations, maintaining a fractional assignment $\bar{x}$ of items to agents; $\bar{x}=x$ initially. Let $\Delta$ be a large enough integer so that every $y_{i,S}$ is an integer multiply of $1/\Delta$. Focus on a fixed agent $i\in{\mathcal{A}}$ and consider the Nash Social Welfare instance containing $\Delta$ copies of $i$, and $\Delta\bar{x}_{ij}$ copies of each item $j\in{\mathcal{G}}$. Group the items as follows: the $\Delta$ most valuable items belong to the first group, the next $\Delta$ most valuable items belong to the second group, and so on. We define $\Phi_{i}$ to be the logarithm of the Nash Social Welfare of the worst allocation satisfying the following condition: every agent gets at most one item from each group. Fortunately, the worst allocation can be defined naturally: the first agent takes the most valuable item from each group, and the second agent takes the second most valuable item from each group, and so on. Thus $\Phi_{i}$ can be computed efficiently. We define $\Phi=\sum_{i}w_{i}\Phi_{i}$ to be the overall potential function. In the randomized version of the algorithm, one can define the rotation operation over the fractional matching between groups $G$ and items ${\mathcal{G}}$. In expectation the operation does not decrease $\Phi$. To derandomize the algorithm, we can perform the operation deterministically so that $\Phi$ does not decrease. The potential value $\Phi$ at the end of the algorithm is at least that at the beginning, which is at least the value of the configuration LP minus $1/e$. On the other hand, the logarithm of the Nash Social Welfare of the integral solution is exactly the final $\Phi$. Therefore, the Nash Social Welfare is at least $e^{-1/e}$ times the exponential of the value of the configuration LP.
Appendix A Solving Configuration LP within an Additive Error of $\ln(1+\epsilon)$
Let $\epsilon>0$ be upper bounded by a sufficiently small constant (we allow $\epsilon$ to be a sub-constant). By only allowing every agent to get one item, we can obtain an $m$-approximation for the our Nash Social Welfare instance. Then, by making $O\Big{(}\frac{\log m}{\epsilon}\Big{)}$ guesses, we can assume we are given a number $o$ such that the value of (Conf-LP) is in $(o,o+\epsilon/3]$.
We consider the dual of (Conf-LP), with the objective replaced by a constraint.
$\sum_{j\in{\mathcal{G}}}\alpha_{j}+\sum_{i\in{\mathcal{A}}}\beta_{i}$ $\leq o$ (4)
$\sum_{j\in S}\alpha_{j}+\beta_{i}$ $\geq w_{i}\cdot\ln v_{i}(S)$ $\forall i\in{\mathcal{A}},S\subseteq{\mathcal{G}}$ (5)
$\alpha_{j}$ $\geq 0$ $\forall j\in{\mathcal{G}}$ (6)
Since (Conf-LP) has value strictly larger than $o$, the dual LP (4-6) is infeasible. We design an approximate separation oracle for the LP. Given some $\alpha\in{\mathbb{R}}_{\geq 0}^{\mathcal{G}}$ and $\beta\in{\mathbb{R}}^{\mathcal{A}}$ that does not satisfy (5), we can find some $i\in{\mathcal{A}}$ and $S\subseteq{\mathcal{G}}$ such that
$\sum_{j\in S}\alpha_{j}+\beta_{i}<w_{i}\ln\big{(}(1+\epsilon/2)v_{i}(S)\big{)}.$
The running time of the oracle is polynomial in the input size and $\frac{1}{\epsilon}$. This can be achieved using the standard dynamic programming technique: For a fixed $i\in{\mathcal{A}}$, to find the $S$, we guess the item $j^{*}\in S$ with the largest $v_{ij^{*}}$, coarsen the $v_{ij}$ values based on the guess, and run a dynamic programming to find the $S$.
So, using the ellipsoid method with the approximate separation oracle, we can find polynomially many half spaces of the form $\sum_{j\in S}\alpha_{j}+\beta_{i}\geq w_{i}\ln\big{(}(1+\epsilon/2)v_{i}(S)\big{)}$, whose intersection is empty. Then, we consider the Nash Social Welfare instance where all $v_{ij}$ values are scaled up by $21+\epsilon/$ 2, and (Conf-LP) to the instance. By solving the LP restricted to the variables $y_{i,S}$ correspondent to the half spaces (that is, we let all other variables be 00), we obtain a solution $y$ whose value is at least $o$ w.r.t the scaled instance. So, the value of the solution $y$ to (Conf-LP) w.r.t the original instance is at least $o-\sum_{i\in{\mathcal{A}},S\subseteq{\mathcal{G}}}y_{i,S}w_{i}\ln(1+\epsilon/2)=o-\sum_{i}w_{i}\ln(1+\epsilon/2)=o-\ln(1+\epsilon/2)$.
As the value of (Conf-LP) is at most $o+\epsilon/3$, we solved the LP up to an additive error of $\epsilon/3+\ln(1+\epsilon/2)$. For a small enough $\epsilon$, this is at most $\ln(1+\epsilon)$.